The sum of May's age and June's age is 80. Ten years ago May was thrice as old as June then. How old is June?
Answer by TimothyLamb(4379)

x = may age now
y = june age now
---
x + y = 80
x - 10 = 3(y - 10)
---
put the system of linear equations into standard form
---
x + y = 80
x - 3y = -20
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = may age now = 55
y = june age now = 25
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---